On a railway track,a driver brakes at a yellow signal at a distance of 1km from red signal, where it stops...the maximum deceleration of train is 0.2m/s square... assuming uniform deceleration,what is maximum safe speed of the train at yellow signal?
1
Expert's answer
2020-04-27T10:14:23-0400

applying formula "v^2-u^2=2aS"

here a=-0.2 m/s2 , v=0 m/s , S=1000 m and u=?

substituting in the formula we get

"0-u^2=2\times(-0.2)\times1000=400m/s"

u=20m/s

The maximum speed the train can have is 20m/s